OpenStudy (anonymous):

Use the quadratic formula to find the zeros of the function y = 2x2 + 3x + 6. Round to the nearest tenth if necessary. A. {-1.7, 2.3} B. {-2.4, 2.4} C. {-2.3, 0.8} D. no real solution

OpenStudy (anonymous):

a=2; b=3; c=6 formula is \[-b \pm \sqrt{b ^{2}-4ac}\]/2a solve!!

OpenStudy (anonymous):

\[-b \pm \sqrt{b ^{2}-4ac}/2a\]

OpenStudy (anonymous):

\[\LARGE {x_{1/2}} = \frac{{ - b \pm \sqrt {{b^2} - 4ac} }}{{2a}}\]

OpenStudy (anonymous):

so for this one: \[-3\pm \sqrt{-39}/4\] there is no real solution b/c you cannot solve the sq. root of -39!!

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

There is no actual solution